There are two octahedral dice with faces marked 1 through 8. They are constructed such that the side 8 is 1.5 times as probable as the 2 through 7 and the sum of the probabilities of the 1 and the 8 equals that of the pairs on opposing faces. Let x be the sum of the faces when we roll the two dice.



Find the following:



a. Probability distribution of x



b. Mean of x



c. Variance of x

1
Expert's answer
2022-02-09T13:47:58-0500

Let the probability of "1" be "y." Let the probability of "2" be "s." Then


"y+s+s+s+s+s+s+1.5s=1""y+1.5s=s+s"

"y+7.5s=1""y=0.5s"

"y=1\/16, s=1\/8"

There are "8^2=64" possible outcomes for "X"


"1+1=2"

"1+2=2+1=3"

"1+3=2+2=3+1=4"


"1+4=2+3=3+2=4+1=5"

"1+5=2+4=3+3=4+2=5+1=6"

"1+6=2+5=3+4=4+3=5+2=6+1=7"

"1+7=2+6=3+5=4+4=5+3=6+2""=7+1=8"

"1+8=2+7=3+6=4+5=5+4=6+3""=7+2=8+1=9"

"2+8=3+7=4+6=5+5=6+4=7+3""=8+2=10"

"3+8=4+7=5+6=6+5=7+4""=8+3=11"


"4+8=5+7=6+6=7+5=8+4=12"

"5+8=6+7=7+6=8+5=13"

"6+8=7+7=8+6=14"

"7+8=8+7=15"


"8+8=16"

"P(X=2)=\dfrac{1}{16}(\dfrac{1}{16})=\dfrac{1}{256}"


"P(X=3)=\dfrac{1}{16}(\dfrac{1}{8})+\dfrac{1}{8}(\dfrac{1}{16})=\dfrac{1}{64}"

"P(X=4)=\dfrac{1}{16}(\dfrac{1}{8})+\dfrac{1}{8}(\dfrac{1}{8})+\dfrac{1}{8}(\dfrac{1}{16})=\dfrac{1}{32}"

"P(X=5)=\dfrac{1}{16}(\dfrac{1}{8})+2(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{1}{8}(\dfrac{1}{16})=\dfrac{3}{64}"

"P(X=6)=\dfrac{1}{16}(\dfrac{1}{8})+3(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{1}{8}(\dfrac{1}{16})=\dfrac{1}{16}"

"P(X=7)=\dfrac{1}{16}(\dfrac{1}{8})+4(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{1}{8}(\dfrac{1}{16})=\dfrac{5}{64}"

"P(X=8)=\dfrac{1}{16}(\dfrac{1}{8})+5(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{1}{8}(\dfrac{1}{16})=\dfrac{3}{32}"

"P(X=9)=\dfrac{1}{16}(\dfrac{3}{16})+6(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{3}{16}(\dfrac{1}{16})=\dfrac{15}{128}"

"P(X=10)=\dfrac{1}{8}(\dfrac{3}{16})+5(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{3}{16}(\dfrac{1}{8})=\dfrac{1}{8}"

"P(X=11)=\dfrac{1}{8}(\dfrac{3}{16})+4(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{3}{16}(\dfrac{1}{8})=\dfrac{7}{64}"

"P(X=12)=\dfrac{1}{8}(\dfrac{3}{16})+3(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{3}{16}(\dfrac{1}{8})=\dfrac{3}{32}"

"P(X=13)=\dfrac{1}{8}(\dfrac{3}{16})+2(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{3}{16}(\dfrac{1}{8})=\dfrac{5}{64}"

"P(X=14)=\dfrac{1}{8}(\dfrac{3}{16})+(\dfrac{1}{8})(\dfrac{1}{8})+\dfrac{3}{16}(\dfrac{1}{8})=\dfrac{1}{16}"

"P(X=15)=\dfrac{1}{8}(\dfrac{3}{16})+\dfrac{3}{16}(\dfrac{1}{8})=\dfrac{3}{64}"

"P(X=16)=\dfrac{3}{16}(\dfrac{3}{16})=\dfrac{9}{256}"

a. Probability distribution of x

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & p(x) \\ \hline
 2& 1/256\\
 \hdashline
 3& 1/64\\
 \hdashline
 4& 1/32\\
 \hdashline
 5& 3/64\\
 \hdashline
6& 1/16\\
 \hdashline
 7& 5/64\\
 \hdashline
 8& 3/32\\
 \hdashline
 9& 15/128\\
 \hdashline
10 &1/8\\
 \hdashline
 11& 7/64\\
 \hdashline
12& 3/32\\
 \hdashline
 13& 5/64\\
 \hdashline
 14& 1/16\\
 \hdashline
15& 3/64\\
 \hdashline
16& 9/256\\
 \hdashline
\end{array}"


b.

"mean=E(X)=2(\dfrac{1}{256})+3(\dfrac{1}{64})+4(\dfrac{1}{32})"

"+5(\dfrac{3}{64})+6(\dfrac{1}{16})+7(\dfrac{5}{64})+8(\dfrac{3}{32})"

"+9(\dfrac{15}{128})+10(\dfrac{1}{8})+11(\dfrac{7}{64})+12(\dfrac{3}{32})"

"+13(\dfrac{5}{64})+14(\dfrac{1}{16})+15(\dfrac{3}{64})+16(\dfrac{1}{256})"

"=9.375"

c. Variance of x


"mean=E(X^2 )=2^2(\dfrac{1}{256})+3^2(\dfrac{1}{64})+4^2(\dfrac{1}{32})"

"+5^2(\dfrac{3}{64})+6^2(\dfrac{1}{16})+7^2(\dfrac{5}{64})+8^2(\dfrac{3}{32})"

"+9^2(\dfrac{15}{128})+10^2(\dfrac{1}{8})+11^2(\dfrac{7}{64})+12^2(\dfrac{3}{32})"

"+13^2(\dfrac{5}{64})+14^2(\dfrac{1}{16})+15^2(\dfrac{3}{64})+16^2(\dfrac{1}{256})"

"=99.6328125"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2"

"=99.6328125-(9.375)^2=11.7421875"